The corner section of a football stadium has 15 seats in the first row and 40 rows in all. Each sucessive row contains 2 additional seats. Howmany total seats are in this section?
Answer by stanbon(75887)

The corner section of a football stadium has 15 seats in the first row and 40 rows in all. Each sucessive row contains 2 additional seats. Howmany total seats are in this section?
------------
The problem involves the sum of an arithmetic series where d=2, a(1)=15
You need to find a(40)=a(1)+39*2=15+78=93
Then you need to find S(40)=(40/2)(93)
=20(93)=1860
Number of seats is 1860
Cheers,
Stan H.